OpenStudy (anonymous):

Gillian purchased 25 books at the library book sale. Each hardcover book cost $1.50, and each paperback book cost $0.50. Gillian spent a total of $26.50. The book costs can be represented by the system of equations below. h + p = 25 1.50h + 0.50p = 26.50 How many paperback books did Gillian buy?

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (secret-ninja):

Sorry, but I have to get started on a Biology project and I wont have time to help! I just wanted to let you know. Have an awesome day. :D

OpenStudy (anonymous):

Thanks

OpenStudy (cj49):

multiply the first equation by 1.5 (h+p=25)

OpenStudy (cj49):

(subtract equation 1 from 2) 1.5h+1.5p=37.5 1.5h+0.5p=26.5 1p=11 p=11

OpenStudy (anonymous):

So the answer is 11 paperback books?

OpenStudy (anonymous):

Or is it 14 ?

OpenStudy (cj49):

yes its 11

OpenStudy (anonymous):

thank you

OpenStudy (cj49):

yw